) Write the equation y=x^2-8x in the form y=a(x-h)^2 +k
Answer by Alan3354(69443)

Write the equation y=x^2-8x in the form y=a(x-h)^2 +k
-----------
y=x^2-8x
complete the square of x^2-8x
x^2-8x+16
--> y+16 = x^2-8x+16
y+16 = (x-4)^2
y = (x-4)^2 - 16
===================
a = 1